# Impact of Charge Transfer Inefficiency on transit light-curves

### v1

@@ 3.6 Polynomial coefficients from theory and fit
where $a_{0}$ represents the constant term and $a_{2}$ the quadratic term. The linear term $a_{1}$ is zero by definition. We provide the theoretical values for $a_{0}$, $a_{1}$, and $a_{2}$ coefficients in Table 4 alongside their fitted counterparts.
Table 4: Fitted and theoretical polynomial coefficients ($a_{0},a_{1},a_{2}$) per species, along with the true values of $\beta$ and $\tau_{\mathrm{r},k}$ from Prod’homme et al. (2016) (see Table 1). We also report the fitted values of $\beta$ and $\tau_{\mathrm{r},k}$, together with their respective $1\sigma$ uncertainties. The release time for the first species was fixed to $\tau_{\mathrm{r},0}=3.05\times 10^{-5}\,\mathrm{s}$ (from Step 1; Table 3), as it is shorter than the line transfer time ($\sim 900\penalty 10000\ \mu\mathrm{s}$) and cannot be constrained by the fit.
@@
From the fitted polynomial coefficients obtained in Step 3 (Sect. 3.5), we derived the cumulative trap density at the top of the CCD, i.e. the quantity $\tilde{\mathrm{cnt}}_{k}[4509,j]$ for columns indexed from $j=0$ to $j=4059$. This estimate can then be directly compared to the simulated (or true reference) cumulative trap density defined by Eq. 7. Fig. 11 illustrates these two cases overplotted for all four species. The result for the first species is poorly constrained, while the second species is mildly constrained and shows a slight systematic offset relative to the true profile. However, as noted previously in Sect. 2.5, the first two species have negligible impact on the CTI-induced bias. Accordingly, the discrepancies found for the first and second species have no consequences.
Ultimately, these calibrated coefficients, capturing the spatial dependence of the trap density, form the basis for the CTI correction procedure described in the next section.
Figure 11: Comparison of cumulative trap density ($\mathrm{cnt}$) at the top of CCD for all species from two sources: the simulation ground truth (in blue), and the fitted polynomial model (in dashed red). The plot highlights certain species-specific behaviour, with Species 1 being poorly constrained and Species 2 exhibiting a systematic offset.
@@ 4 CTI correction strategy
In this section, we describe a strategy to correct CTI in imagettes acquired with the normal cameras at a 25 s cadence. The correction builds on the algorithm of Massey et al. (2014) and uses our calibrated trap-density map.
@@ 4.1 Correction procedure
The fitted polynomial coefficients obtained in the previous section provided the basis for constructing the trap density map necessary for CTI correction of the planetary transit discussed in Sect 2.5. Using Eq. 15, we defined the per-pixel trap densities across the full CCD area as
@@
where $a_{0,k}^{\mathrm{fit}}$, $a_{1,k}^{\mathrm{fit}}$, and $a_{2,k}^{\mathrm{fit}}$ are the fitted polynomial coefficients for species $k$ (see Table 4), and $R_{i,j}$ is the normalised radial coordinate at pixel $(i,j)$ as defined in Eq. 2.
The cumulative number of traps per pixel is computed by summing $n_{t,k}[i,j]$ along the CCD rows (see Eq. 13). This cumulative trap density map forms the basis for correcting CTI-affected planetary transits within the imagette containing the source.
Imagettes acquired during observations are small ($\times6$ pixels), but charges deferred by CTI from sources located below the imagette can enter its area, introducing contamination. The CTI effect on a given target also depends on the filling history of the traps, which is determined by the charge distribution in pixels located below the imagette along the readout direction. To account for this, we shall use the full-frame images acquired at the start of each quarter, which provide the flux distribution outside the imagettes.
The CTI correction is implemented using an adaptation of the iterative scheme proposed by Massey et al. (2014), applied to each imagette. The algorithm employed the routine run_cdm_parallel_cumul, using the cumulative number of traps, $\mathrm{cnt}_{k}[i,j]$ as input (together with all the other constant CTI parameters). Each column is iteratively corrected for a maximum of 10 iterations or until the reduced $\chi^{2}$ fell below a threshold of $0.001$. After correction, photometry was extracted from the imagettes using an optimal aperture mask, following the procedure described in Sect. 2.5.
@@ 4.2 Correction results and performance evaluation
The final transit depths (in ppm) were computed for each case. We defined the residual correction below
@@
where $D_{\mathrm{no\ CTI}}$ denotes the transit depth without CTI, and $D_{\mathrm{corr}}$ denotes the transit depth after CTI correction.
This framework enabled us to assess the performance of the CTI correction strategy in recovering the true transit depth and, more broadly, the overall efficiency of the method.
We summarize the CTI-induced bias, the residuals after correction, and the corresponding improvement in Table 5. Results are presented for a star with $m_{V}=11$ at positions A, B, and C (see Fig. 4) for mission durations of 8 and 4 years. For completeness, we also include two extreme cases: a bright star ($m_{V}=9$) and a faint star ($m_{V}=13$), both at position C under the 8-year scenario, which represents the worst-case CTI impact.
Table 5: CTI-induced bias and correction performance for different mission durations, positions, and magnitudes. Bias represents the CTI impact, Residual shows the remaining error after correction, and Improvement quantifies the fraction of bias removed.
@@
Fig. 12 compares the pristine, CTI-affected, and CTI-corrected transit depths for a star with $m_{V}=11$, representing the worst-case scenario (8-year mission duration, position C). The initial CTI bias of $3.95\%$ was reduced to a residual of $0.06\%$, corresponding to an improvement of $98.61\%$, well within PLATO’s compliance window.
@@
Figure 12: Comparison of transit light-curves for the worst-case scenario: an 8-year mission, high trap density (position C), and $m_{V}=11$. The solid orange curve corresponds to the CTI-affected case, the solid green curve to the CTI-corrected case (nearly overlapping the dashed blue curve representing the CTI-free case). For an initial bias of $3.95\%$, the correction reduces the residual to $0.06\%$, achieving an improvement of $98.61\%$. The lower panel shows the correction residual.
@@ 5.1 Conclusions
In response to this challenge, we designed a calibration and correction strategy for CTI. A dedicated calibration dataset was generated using PLATOSim (Jannsen et al., 2024), simulating a realistic stellar distribution over the full CCD. To capture spatial variations in the trap density, we model its axisymmetric radial dependence using a polynomial function of the distance from the optical centre. The calibration begins with an empirical fit to determine the number of trap species and their initial release time estimates $\tau_{\mathrm{r},k}$. We then model and remove the smearing contribution, which would otherwise bias subsequent steps. Next, an iterative algorithm adapted from Massey et al. (2014), combined with the EPER method (Prod’homme et al., 2016), is applied to constrain, for each species, the polynomial coefficients $a_{p,k}$ describing the spatial variation of trap density, as well as the well-fill power-law index $\beta$ and the release time constants $\tau_{\mathrm{r},k}$.
Finally, the fitted coefficients are used in a column-wise iterative correction procedure, also adapted from Massey et al. (2014), to approximate CTI-free images. This correction reduces the systematic bias in transit depth measurements to a level fully compliant with the mission’s requirements.
@@ 5.2 Discussion and perspectives
Our calibration currently assumes access to the full image, including regions below the imagette containing the target. In practice, this assumption only holds at the beginning of a quarter, since full-frame images will not be available during the observations. Consequently, contaminant stars located below the imagette will drift over time, complicating CTI correction. Another limitation arises from our reliance on aperture photometry, which is not optimized for crowded stellar fields where close companions cannot be disentangled. A PSF-fitting approach is implemented in the PLATO pipeline to overcome these issues. Assessing the CTI correction efficiency with such photometry extraction method is an important future step (work in progress). For light-curves generated on-board using aperture photometry, the CTI correction strategy proposed in this work cannot be applied in-flight. For these cases, a dedicated on-ground correction algorithm will be required, likely leveraging knowledge of the PSF, since pixel-level information is unavailable for stars whose photometry is computed on-board.
@@
In our simulation, the long-term drift of the star along the focal plane does not affect the transit depth inference. This is because the simulated light-curves have been de-trended prior to measuring the transit depth. The PLATO light-curves will be detrended using a variant of the PCD-MAP algorithm (Smith et al., 2012), optimized for PLATO. However, this de-trending will not be perfect and may leave residuals that could introduce bias in the transit depth measurement.
@@
Our model for the spatial variation in trap density assumes a radial distribution with a gradual increase toward the edges of the focal plane. This assumption is primarily supported by radiation analysis for PLATO cameras, although the analysis reveals moderate large-scale departures from perfect axisymmetry. Improving this model will require introducing azimuthal components, which remain to be defined. In principle, analyzing CTI-induced trails from stellar images at different CCD positions could refine the model and enhance correction accuracy; however, this approach warrants a dedicated feasibility study.
@@
The capture cross-sections $\sigma_{k}$ for each species were fixed using the values reported in Prod’homme et al. (2016), which are consistent with those found in the literature (Prod’homme et al., 2014a; Hall et al., 2015). The EPER method, however, does not allow for the derivation of capture cross-sections. To obtain these values, one must instead use the First Pixel Response (FPR) method (see e.g. Prod’homme et al., 2014a, b). The FPR method involves studying the charge loss that occurs at the leading edge of the charge block transferred through the CCD due to CTI. However, it remains an open question whether PLATO data can be used effectively to apply the FPR method, which is a topic we plan to address in the future.
@@
We assume that the radiation dose increases slowly over time, with calibration planned at the start of each three-month observation period. This assumption may fail in the event of major solar activity, such as flares, which can cause rapid surges in radiation. In such cases, additional calibration using CCD images acquired immediately after the event would be necessary.
@@
Our validation is based entirely on simulated datasets, as it is infeasible to corroborate the results with real in-flight images. A promising path forward is to test the correction scheme using transit light-curves obtained during operations with different cameras. Indeed, CTI will not affect all CCDs and cameras uniformly. As shown in Fig. 13, Camera 9 exhibits the lowest radiation exposure, in contrast to Camera 1 in Fig. 2. Such variations, both across cameras and among CCDs within a camera, translate into differences in trap distribution and, consequently, in the severity of CTI. Hence, comparing transit light-curves obtained on cameras with very different radiation damage should in principle allow us to validate the efficiency of the CTI correction.
@@
Figure 13: Radiation map of camera 9 illustrating its lower radiation exposure compared to camera 1 in Fig. 2, as camera 9 is located near the centre of the optical bench whereas camera 1 is positioned at a corner. Adapted from OHB System AG (2024).
@@
Addressing all these challenges, in parallel with refining our correction algorithm, is essential to ensure PLATO meets its scientific objectives in the detection and characterization of Earth-sized exoplanets. These challenges will be addressed in future work.
@@ Acknowledgements.
Appendix A Python implementation
import numpy as np
from numba import prange,njit
@njit(nogil=True,fastmath=True,cache=True,parallel=True)
def run_cdm_parallel_cumul(
image: np.ndarray,
y0: int,
beta: float,
vg: float,
t: float,
fwc: float,
vth: float,
tr: np.ndarray,
cnt: np.ndarray,
sigma: np.ndarray,
ci: bool = False
) -> np.ndarray:
r"""
Run ‘CDM‘ in the parallel direction.
This version was adapted from the routine run_cdm_parallel() taken from Pyxel 1.10.2
(https://esa.gitlab.io/pyxel/).
It accounts for the spatial variation of the traps across the CCD.
For more details about Pyxel see:
Arko, M. et al. 2022, in Society of Photo-OpticalInstrumentation Engineers (SPIE) Conference Series,
ol. 12187, Modeling,Systems Engineering, and Project Management for Astronomy X,
ed. G. Z. 660Angeli & P. Dierickx, 1218705
Parameters
image: ndarray
Input image.
y0: int
location of the first row in the CCD (for instance y0=2000 means that the first row in the
input array corresponds to CCD row of index 2000)
beta: float
[…]
vg: float
Maximum geometrical volume :math:‘V_g‘ that electrons can occupy within a pixel. Unit: :math:‘cm^3‘.
t: float
Transfer period :math:‘t‘ (TDI period). Unit: :math:‘s‘.
fwc: float
Full well capacity :math:‘FWC‘. Unit: :math:‘e^-‘.
vth: float
Electron thermal velocity.
tr: sequence of float
[…]
cnt[4510,nx,ns] : cumulated number of traps along the column (i) up to row (j) for the species (k)
sigma: sequence of float
[…]
ci: bool
frame, the input frame assumed to contain CI and it is transferred into the CCD area.
Default: False
Returns
array: ndarray
Output array.
array = image.copy() # we do a copy to preserve the input image
ydim, xdim = array.shape # full signal array we want to apply cdm for
kdim_p = cnt.shape[2]
# IMAGING (non-TDI) MODE
# Parallel direction
no = np.zeros((xdim, kdim_p))
alpha_p: np.ndarray = t * sigma * vth * fwc**beta / (2.0 * vg)
g_p: np.ndarray = 2.0 * cnt / fwc**beta
for j in prange(xdim):
for i in range(ydim):
for k in range(kdim_p):
nc = 0.0
if array[i,j] > 0.01:
if (ci):
if (i+y0 < 4510):
gamma = g_p[4509,j,k] - g_p[i+y0,j,k]
else:
gamma = 0.
else:
if (i+y0 < 4510):
gamma = g_p[i+y0,j,k]
else:
gamma = g_p[4509,j,k]
nc = max(
( gamma*array[i,j] ** beta - no[j, k])
/ ( gamma*array[i,j] ** (beta - 1.0) + 1.0)
* (1.0 - np.exp(-1 * alpha_p[k] * array[i,j] ** (1.0 - beta))),
0.0) # Eq. 22 in Short et al (2013)
no[j,k] += nc
nr = no[j,k] * (1.0 - np.exp(-t / tr[k]))
array[i,j] += -1 * nc + nr
no[j,k] -= nr
if array[i,j] < 0.01:
array[i,j] = 0.0
return array
@njit(nogil=True,fastmath=True,cache=True,parallel=True)
def run_cdm_parallel_cumul_radial_poly(
image: np.ndarray,
colindex: np.ndarray,
y0: int,
beta: float,
vg: float,
t: float,
fwc: float,
vth: float,
tr: np.ndarray,
a: np.ndarray,
xc: float,
yc: float,
sigma: np.ndarray,
ci: bool = False
) -> np.ndarray:
r"""
Run :term:‘CDM‘ in parallel direction while taking account the trap distribution across the CCD.
Here the trap density is modelled on the basis of a polynomial model
varying as a function of the distance from the centre of the focal plane.
Adapted from run_cdm_parallel() taken from Pyxel 1.10 (https://esa.gitlab.io/pyxel/).
Parameters
image: ndarray
Input image.
colindex: integer array
index of the columns provided in the input image
y0: int
location of the first row in the CCD covered by input array (for instance y0=2000 means that the
first row in the input array corresponds to CCD row of index 2000)
beta: float
[…]
vg: float
Maximum geometrical volume :math:‘V_g‘ that electrons can occupy within a pixel. Unit: :math:‘cm^3‘.
t: float
Transfer period :math:‘t‘ (TDI period). Unit: :math:‘s‘.
fwc: float
Full well capacity :math:‘FWC‘. Unit: :math:‘e^-‘.
vth: float
Electron thermal velocity.
tr: sequence of float
[…]
a: ndarray (ns,deg+1)
Polynomial coefficients modelling the trap density as a function of the radius (from the
optical centre, ns is the number of species and deg the degree of the polynomial function
[…]
xc,yc: (int,int)
positions of the principal point (in the CCD reference frame)
sigma: sequence of float
[…]
ci: bool
frame, the input frame assumed to contain CI and it is transferred into the CCD area.
Default: False
Returns
array: ndarray
Output array.
array = image.copy() # we do a copy to preserve the input image
ydim, xdim = array.shape # full signal array we want to apply cdm for
kdim_p = a.shape[0] # number of species
# IMAGING (non-TDI) MODE
# Parallel direction
no = np.zeros((xdim, kdim_p))
alpha_p: np.ndarray = t * sigma * vth * fwc**beta / (2.0 * vg)
#< derive the cumulated trap density (not computed for all the CCD rows)
cnt = np.zeros((kdim_p,4510,xdim)) # cumulated trap density
Np = a.shape[1]
D = np.zeros((4510,xdim,Np))
for i in prange(y0,4510):
for j in range(xdim):
RsqX = ((colindex[j] - xc) / 4510) ** 2 / 2.
for p in range(Np):
for ip in range(i+1):
R = math.sqrt(RsqX + ((ip - yc) / 4510) ** 2 /2.)
D[i,j,p] += R**p
for i in prange(y0,4510):
for k in range(kdim_p):
for j in range(xdim):
for p in range(Np):
cnt[k,i,j] += a[k,p]*D[i,j,p]
g_p: np.ndarray = 2.0 * cnt / fwc**beta
for j in prange(xdim):
for i in range(ydim):
for k in range(kdim_p):
nc = 0.0
if array[i,j] > 0.01:
if (ci):
if (i+y0 < 4510):
gamma = g_p[k,4509,j] - g_p[k,i+y0,j]
else:
gamma = 0.
else:
if (i+y0 < 4510):
gamma = g_p[k,i+y0,j]
else:
gamma = g_p[k,4509,j]
nc = max(
( gamma*array[i,j] ** beta - no[j,k])
/ ( gamma*array[i,j] ** (beta - 1.0) + 1.0)
* (1.0 - np.exp(-1 * alpha_p[k] * array[i,j] ** (1.0 - beta))),
0.0) # Eq. 22 in Short et al (2013)
no[j,k] += nc
nr = no[j,k] * (1.0 - np.exp(-t/ tr[k]))
array[i,j] += -1 * nc + nr
no[j,k] -= nr
if array[i,j] < 0.01:
array[i,j] = 0.0
return array

### v2

@@ 3.6 Polynomial coefficients from theory and fit
where $a_{0}$ represents the constant term and $a_{2}$ the quadratic term. The linear term $a_{1}$ is zero by definition. We provide the theoretical values for the $a_{0}$, $a_{1}$, and $a_{2}$ coefficients in Table 4 alongside their fitted counterparts.
Table 4: Fitted and theoretical polynomial coefficients ($a_{0},a_{1}$, and $a_{2}$) per species. [^7] This table includes the true values of $\beta$ and $\tau_{\mathrm{r},k}$ from Prod’homme et al. (2016, see our Table 1) and reports the fitted values together with their respective $1\sigma$ uncertainties. The release time for the first species was fixed to $\tau_{\mathrm{r},0}=3.10\times 10^{-5}\,\mathrm{s}$ (from Step 1; Table 3), as it is shorter than the line transfer time ($\sim 900\penalty 10000\ \mu\mathrm{s}$) and cannot be constrained by the fit.
@@
From the fitted polynomial coefficients obtained in Step 3 (Sect. 3.5), we derived the cumulative trap density at the top of the CCD, i.e. the quantity $\tilde{\mathrm{cnt}}_{k}[4509,j]$ for columns indexed from $j=0$ to $j=4059$. This estimate can then be directly compared to the simulated (or true reference) cumulative trap density defined by Eq. 7.
Figure 11 illustrates these two cases overplotted for all four species. The result for the first species is poorly constrained, while the second species is mildly constrained and shows a slight systematic offset relative to the true profile. However, as noted previously in Sect. 2.5, the first two species have negligible impact on the CTI-induced bias. Accordingly, the discrepancies found for the first and second species have no consequences. Ultimately, these calibrated coefficients, capturing the spatial dependence of the trap density, form the basis for the CTI correction procedure described in the next section.
Figure 11: Comparison of cumulative trap density ($\mathrm{cnt}$) at the top of the CCD for all species from two sources: the simulation ground truth (in blue) and the fitted polynomial model (in dashed red). The plot highlights certain species-specific behaviour, with Species 2 exhibiting a systematic offset. Note that Species 1 is poorly constrained because its release time is significantly shorter than the line transfer time; however, as detailed in Sect. 2.5, this species contributes negligibly to the total CTI-induced bias.
@@ 4 CTI correction strategy
In this section we describe a strategy to correct CTI in imagettes acquired with the normal cameras at a 25 s cadence. The correction builds on the algorithm of Massey et al. (2014) and uses our calibrated trap-density map.
@@ 4.1 Correction procedure
The fitted polynomial coefficients obtained in the previous section provided the basis for constructing the trap density map necessary for CTI correction of the planetary transit discussed in Sect. 2.5. Using Eq. 15, we defined the per-pixel trap densities across the full CCD area as
@@
where $a_{0,k}^{\mathrm{fit}}$, $a_{1,k}^{\mathrm{fit}}$, and $a_{2,k}^{\mathrm{fit}}$ are the fitted polynomial coefficients for species $k$ (see Table 4), and $R_{i,j}$ is the normalised radial coordinate at pixel $(i,j)$ as defined in Eq. 2. The cumulative number of traps per pixel is computed by summing $n_{t,k}[i,j]$ along the CCD rows (see Eq. 13). This cumulative trap density map forms the basis for correcting CTI-affected planetary transits within the imagette containing the source.
Imagettes acquired during observations are small ($\times6$ pixels), but charges deferred by CTI from sources read out before the imagettes can enter its area, introducing contamination. In order to take this into account, it is necessary to know the filling history of the traps, which is determined by the charge distribution in pixels read out before the imagettes along the readout direction. To account for this, we used the full-frame images acquired at the start of each quarter, which provide the flux distribution outside the imagettes.
The CTI correction is implemented using an adaptation of the iterative scheme proposed by Massey et al. (2014), applied to each imagette. The algorithm employed the routine run_cdm_parallel_cumul, using the cumulative number of traps ($\mathrm{cnt}_{k}[i,j]$) as input (together with all the other constant CTI parameters). Each column is iteratively corrected for a maximum of 10 iterations or until the reduced $\chi^{2}$ fell below a threshold of $0.001$. After correction, photometry was extracted from the imagettes using an optimal aperture mask, following the procedure described in Sect. 2.5.
@@ 4.2 Correction results and performance evaluation
The final transit depths (in ppm) were computed for each case. We defined the residual correction as
@@
where $D_{\mathrm{no\ CTI}}$ denotes the transit depth without CTI, and $D_{\mathrm{corr}}$ denotes the transit depth after CTI correction. This framework enabled us to assess the performance of the CTI correction strategy in recovering the true transit depth and, more broadly, the overall efficiency of the method. We summarise the CTI-induced bias, the residuals after correction, and the corresponding improvement in Table 5. Results are presented for a star with $m_{V}=11$ at positions A, B, and C (see Fig. 4) for mission durations of 8 and 4 years. For completeness, we also include two extreme cases: a bright star ($m_{V}=9$) and a faint star ($m_{V}=13$), both at position C under the 8-year scenario, which represents the worst-case CTI impact.
Table 5: CTI-induced bias and correction performance. [^8] Results are shown for different mission durations, positions, and magnitudes. Bias represents the CTI impact, Residual shows the remaining error after correction, and Improvement quantifies the fraction of bias removed.
@@
Figure 12 compares the pristine, CTI-affected, and CTI-corrected transit depths for a star with $m_{V}=11$, representing the worst-case scenario (8-year mission duration, position C). The initial CTI bias of $3.95\%$ was reduced to a residual of $0.06\%$, corresponding to an improvement of $98.61\%$, well within PLATO’s compliance window.
@@
Figure 12: Top: Comparison of transit light curves for the worst-case scenario: an 8-year mission, high trap density (position C), and $m_{V}=11$. The solid orange curve corresponds to the CTI-affected case, the solid green curve to the CTI-corrected case (nearly overlapping the dashed blue curve representing the CTI-free case). For an initial bias of $3.95\%$, the correction reduces the residual to $0.06\%$, achieving an improvement of $98.61\%$. Bottom: Correction residual.
@@ 5.1 Conclusions
In response to this challenge, we designed a calibration and correction strategy for CTI. A dedicated calibration dataset was generated using PLATOSim (Jannsen et al., 2024), simulating a realistic stellar distribution over the full CCD. To capture spatial variations in the trap density, we modelled its axisymmetric radial dependence using a polynomial function of the distance from the optical centre. The calibration began with an empirical fit to determine the number of trap species and their initial release time estimates ($\tau_{\mathrm{r},k}$). We then modelled and removed the smearing contribution, which would otherwise bias subsequent steps. Next, an iterative algorithm adapted from Massey et al. (2014), combined with the EPER method (Prod’homme et al., 2016), was applied to constrain, for each species, the polynomial coefficients ($a_{p,k}$) describing the spatial variation of trap density, as well as the well-fill power-law index ($\beta$) and the release time constants ($\tau_{\mathrm{r},k}$).
Finally, the fitted coefficients were used in a column-wise iterative correction procedure, also adapted from Massey et al. (2014), to approximate CTI-free images. This correction reduces the systematic bias in transit depth measurements to a level fully compliant with the mission’s requirements.
@@ 5.2 Discussion and perspectives
Our calibration currently assumes access to the full image, including regions read out before the imagette containing the target. In practice, this assumption only holds at the beginning of a quarter, since full-frame images will not be available during the observations. Consequently, sources located in regions read out before the imagettes will drift over time, complicating CTI correction.
@@
Another limitation arises from our reliance on aperture photometry, which is not optimised for crowded stellar fields where close companions cannot be disentangled. A PSF-fitting approach is implemented in the PLATO pipeline to overcome these issues. Assessing the CTI correction efficiency with such a photometry extraction method is an important future step (work in progress). For light curves generated on board using aperture photometry, the CTI correction strategy proposed in this work cannot be applied in-flight. For these cases, a dedicated on-ground correction algorithm will be required, likely leveraging knowledge of the PSF, since pixel-level information is unavailable for stars whose photometry is computed on board.
@@
In our simulation, the long-term drift of the star along the focal plane does not affect the transit depth inference. This is because the simulated light curves will have been de-trended prior to the transit depth being measured. The PLATO light curves will be de-trended using a variant of the PCD-MAP algorithm (Smith et al., 2012) optimised for PLATO.
However, in realistic conditions, a complex coupling between CTI and drift is expected, in contrast to the idealised case of our simulations. Therefore, under actual operations, this de-trending will not be perfect and may leave residuals that could introduce bias in the transit depth measurement.
@@
Our model for the spatial variation in trap density assumes a radial distribution with a gradual increase towards the edges of the focal plane. This assumption is primarily supported by radiation analysis for PLATO cameras, although the analysis reveals moderate large-scale departures from perfect axisymmetry. Improving this model will require introducing azimuthal components, which remain to be defined. In principle, analysing CTI-induced trails from stellar images at different CCD positions could refine the model and enhance correction accuracy; this approach warrants a dedicated feasibility study.
@@
The capture cross-sections ($\sigma_{k}$) for each species were fixed using the values reported in Prod’homme et al. (2016), which are consistent with those found in the literature (Prod’homme et al., 2014a; Hall et al., 2015). The EPER method, however, does not allow for the derivation of capture cross-sections. To obtain these values, one must instead use the first pixel response (FPR) method (see e.g. Prod’homme et al., 2014a, b). The FPR method involves studying the charge loss that occurs at the leading edge of the charge block transferred through the CCD due to CTI. However, it remains an open question whether PLATO data can be used effectively to apply the FPR method, which is a topic we plan to address in the future.
@@
We assume that the radiation dose increases slowly over time, and so calibration is planned at the start of each three-month observation period. This assumption may fail in the event of major solar activity, such as flares, which can cause rapid surges in radiation. In such cases, additional calibration using CCD images acquired immediately after the event would be necessary.
@@
Our validation is based entirely on simulated datasets, as it is infeasible to corroborate the results with real in-flight images. A promising path forwards is to test the correction scheme using transit light curves obtained during operations with different cameras. Indeed, CTI will not affect all CCDs and cameras uniformly. As shown in Fig. 13, Camera 9 exhibits the lowest radiation exposure, in contrast to Camera 1 in Fig. 2. Such variations, both across cameras and among CCDs within a camera, translate into differences in trap distribution and, consequently, in the severity of CTI. Hence, comparing transit light curves obtained on cameras with very different radiation damage should in principle allow us to validate the efficiency of the CTI correction.
@@
Figure 13: Radiation map of Camera 9 illustrating its lower radiation exposure compared to Camera 1 in Fig. 2. Camera 9 is located near the centre of the optical bench, whereas Camera 1 is positioned at a corner. Adapted from OHB System AG (2024).
@@
Addressing all these challenges, in parallel with refining our correction algorithm, is essential to ensuring PLATO meets its scientific objectives in the detection and characterisation of Earth-sized exoplanets. These challenges will be addressed in future work.
Data availability
The CTI calibration and correction pipeline described in this paper is available on GitHub at github.com/mishrashaunak/plato-cti-correction. The underlying data generated for the simulations are available from the corresponding author upon request.